### v1

@@ 4 Conclusions
This paper presents a rapid updating algorithm designed explicitly for multivariate resource models. The algorithm combines EnKF-MDA and RBIG, where the former offers more accurate updates compared to the traditional EnKF, while RBIG transforms multivariate data into independent multi-Gaussian factors that are suitable for updating. This approach is applicable to any mining deposit that has multiple cross-correlated quantitative variables. The effectiveness of the proposed algorithm was demonstrated using real fused data provided by Petra Data Science, achieving an error reduction ranging from 77% to 98% across nine cross-correlated variables. EnKF-MDA enhances the prediction accuracy with five variables, reducing their errors by more than 90%. RBIG satisfies the Gaussian assumption of EnKF-MDA and maintains multivariate relationships.
In addition to the accuracy, the proposed algorithm operates with impressive speed, even on low-cost hardware. For instance, it can update 100 block model realisations with nearly 900 observations from period 1 in under five minutes. The process could be further enhanced through parallelisation across multiple virtual machines, which is easily achievable in an industrial setting. Notably, the rapid updating requires no human intervention and can be automated to run either at specified time intervals or immediately when new information is available. As a result, mining operations can make near real-time decisions based on updated models, leading to improved short-term mine planning and optimisation.
However, some challenges need to be addressed in future research regarding rapid resource model updating. The uncertainty associated with soft data, such as sensor observations, extends beyond measurement errors typically identified in laboratory settings. The way sensors are used on an actual mine site can differ from how they are intended to be used. Not to mention the difference between harsh mining conditions and a laboratory environment where measurement errors are evaluated. Furthermore, although more practical, downstream sensors, such as those on conveyor belts, are hard to link back to resource models. This challenge is particularly relevant in underground mining, where ore tracking remains a significant issue for mining companies. Thus, the uncertainty introduced by ore tracking must be considered during rapid updating.
Quantitative grade variables are not the only critical components of resource models. Qualitative variables, such as lithology, alterations and various domain types, are also essential for resource modelling, optimisation, and mine planning. The primary challenge with qualitative information is that ensemble-based data assimilation mainly operates on Gaussian values. One approach to addressing this challenge is to parametrise qualitative variables into quantitative ones. For example, discrete wavelet transforms can be used to turn geological domains into frequency coefficients before data assimilation (Talesh Hosseini et al, 2023). Alternatively, domains may be represented as multiple Gaussian random fields through pluri-Gaussian simulation (Armstrong et al, 2011).
It is also important to rapidly update geometallurgical models as these parameters can be used to optimise the entire mining value chain. However, this is challenging due to the limited availability of geometallurgical tests for accurate predictions, coupled with many geometallurgical parameters that are non-additive. Nevertheless, applying EnKF to update the Bond Ball Mill Work index helped reduce prediction errors by 72 % and improved future forecasts by 26 % (Wambeke et al, 2018). Future research will focus on expanding the data assimilation algorithm to enable the updating of qualitative variables (e.g., geological domains) and geometallurgical properties.
@@ Acknowledgments
The research reported here was supported by the Australian Research Council Industrial Transformation Training Centre for Integrated Operations for Complex Resources (ARC ITTC IOCR - project number IC190100017) and funded by universities, industry and the Australian Government. We acknowledge Petra Data Science for providing a fused dataset for our experiments.

### v2

@@ 5 Conclusions
This paper presents a rapid updating algorithm designed explicitly for multivariate resource models. The algorithm combines EnKF-MDA and RBIG, where the former offers more accurate updates compared to the traditional EnKF, while RBIG transforms multivariate data into independent multi-Gaussian factors that are suitable for updating. This approach is applicable to any mining deposit that has multiple cross-correlated quantitative variables. The effectiveness of the proposed algorithm was demonstrated through a real case study of an iron ore mine in Western Australia.
Quantitative grade variables are not the only critical components of resource models. Qualitative (e.g., lithology, alterations and various domain types) and non-additive geometallurgical variables (e.g., operating work index, ore recovery, Axb, etc.) are also essential for resource modelling, optimisation, and mine planning. The primary challenge with these types of variables is that ensemble-based data assimilation mainly operates on Gaussian values. However, EnKF has been applied to update geological domains by using discrete wavelet transforms (Talesh Hosseini et al, 2023). Additionally, EnKF has been effective in updating the Bond Ball Mill Work index and improving its future forecasts by 26% (Wambeke et al, 2018).
Future research will focus on expanding the rapid updating algorithm to enable the updating of qualitative variables and geometallurgical properties. The limitations of the current approach will be investigated further, especially the updating of extreme values in highly skewed distributions. Finally, it is important to test how data assimilation performs in a real underground mining scenario and its applicability in enhancing short-term mine planning.
@@ Acknowledgments
The research reported here was supported by the Australian Research Council Industrial Transformation Training Centre for Integrated Operations for Complex Resources (ARC ITTC IOCR - project number IC190100017) and funded by universities, industry and the Australian Government. We acknowledge Petra Data Science for providing a fused dataset for the case study.